1.How many ways can a club consisting of 30 people elect a president, vice president, secretary, and treasurer?
2. Jennifer is asked to complete 5 of the 7 essay questions on her literature esam. How many choices does she have?

Answer by stanbon(75887)

1.How many ways can a club consisting of 30 people elect a president, vice president, secretary, and treasurer?
30 ways to choose the pres
29 ways to choose the vp
28 ways to chose the sec
26 ways to chose the tr
---
Ans: 20P4 = 20!/(20-4)! = 20!/16! = 20*19*18*17 ways
==================================================
2. Jennifer is asked to complete 5 of the 7 essay questions on her literature esam. How many choices does she have?
Ans: 7C5 = 7C2 = (7*6)/(1*2) = 21 choices
=========================================
Cheers,
stan H.
===========